OpenStudy (anonymous):

(x^2+4x-1)/[(x+2)^2]=0 Solve Show steps please.

OpenStudy (nottim):

First, in the second bracket, change it to (x+2)(x+2)

OpenStudy (nottim):

Then, if possible, do trinomials in the first bracket. IF unable, go to the quadratic formula.

OpenStudy (stacey):

In order for a fraction to equal zero, The numerator must equal zero. The denominator can be anything but zero (if it equals zero the fraction is undefined). So to solve this solve x^2 + 4x - 1 = 0.

OpenStudy (anonymous):

x^2 + 4x - 1 = 0 which is?

OpenStudy (anonymous):

use quadratic formula?

OpenStudy (pokemon23):

its mimi

OpenStudy (mimi_x3):

Use the quadratic formula. \[x = \frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2(a)} \] \[a = 1 , b = 4 , c = -1\] And substitute it in.

OpenStudy (mimi_x3):

yes? @ pokemon

OpenStudy (pokemon23):

can you help me does not involve math

OpenStudy (mimi_x3):

What topic ?

OpenStudy (pokemon23):

Health

OpenStudy (mimi_x3):

Health what ? P.D ?

OpenStudy (pokemon23):

like cholesterol

OpenStudy (mimi_x3):

What's the question i'll see if i can help.

OpenStudy (pokemon23):

here a link